# Geometric structure of shallow neural networks and constructive ${\mathcal{L}}^{2}$ cost minimization

### v2

@@ 3.1. Upper bound on minimum of cost function for $M>Q$
is invertible, and we can define the Penrose inverse of $\overline{X_{0}^{red}}$,
${\rm Pen}[\overline{X_{0}^{red}}]:=\big{(}(\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\big{)}^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}$ (3.11)
@@
and we note that ${\rm Pen}[\overline{X_{0}^{red}}]\overline{X_{0}^{red}}={\bf 1}_{Q\times Q}$. The projector property $P^{2}=P$ is thus easily verified, and orthogonality with respect to the Euclidean inner product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on ${\mathbb{R}}^{M}$ holds due to $P^{T}=P$, whereby $(v,Pw)=(Pv,w)$ for any $v,w\in{\mathbb{R}}^{M}$. In particular, we have
@@
Moreover, we define
@@
which measures the relative size between $\overline{X_{0}^{red}}$ and $P\Delta X_{0}$, as $\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|$ scales like the noise to signal ratio $\frac{|\Delta x|}{|x|}$ of training inputs.
@@
Therefore, the upper bound in (3.19) is scaling invariant, as it only depends on the signal to noise ratio of training input data, which is controlled by $\delta_{P}$.
@@ 3.2. Exact degenerate local minimum in the case $M=Q$
We explicitly determine an exact local degenerate minimum of the cost function in the case $Q=M$ where the input and output spaces have the same dimension; here, we show that the upper bound obtained above differs from the sharp value by a relative error of order $O(\delta_{P}^{2})$. We note that in this situation, $P={\bf 1}_{Q\times Q}$ and $P^{\perp}=0$. Moreover, $\overline{X_{0}^{red}}$ is invertible, and therefore, ${\rm Pen}[\overline{X_{0}^{red}}]=(\overline{X_{0}^{red}})^{-1}$.
@@ 3.3. Geometric interpretation
The proof of Theorem 3.1 provides us with a particular set of weights $W_{1}^{*},W_{2}^{*}$ and biases $b_{1}^{*},b_{2}^{*}$ through an explicit construction, which yield the upper bound (3.19). We will refer to the shallow network trained with this choice of $W_{i}^{*},b_{i}^{*}$ as the constructively trained shallow network. A detailed discussion is provided in Section 6.
@@
and matching an input $x\in{\mathbb{R}}^{M}$ with an output $y_{j^{*}}$ via the constructively trained shallow network is equivalent to the solution of the metric minimization problem
@@ 3.4. Dependence on truncation
We define the truncation map $\tau_{W_{1},b_{1}}$ as follows. It will play a major role in our explicit construction of local and global cost minimizers in Deep Learning networks in [4].
@@
Let $W_{1}\in GL(Q)$, $b_{1}\in{\mathbb{R}}^{Q}$. Then, the truncation map $\tau_{W_{1},b_{1}}:{\mathbb{R}}^{Q\times N}\rightarrow{\mathbb{R}}^{Q\times N}$ is defined by
@@
measures the signal to noise ratio of the truncated training input data.
@@
where $P_{R}$ has $Q$ diagonal entries of $1$, and all other components are 00, while $P^{\perp}_{R}$ has $M-Q$ diagonal entries of $1$, and all other components are 00. This does not characterize $R$ uniquely, as we may compose $R$ with an arbitrary $R^{\prime}\in O(M)$ that leaves the ranges of $P,P^{\perp}$ invariant. Diagonality of $P_{R},P_{R}^{\perp}$ implies that the activation function, which acts component-wise as the ramp function (2.9), satisfies
@@
is a matrix all of whose components are $\leq 0$. Therefore, $\sigma$ maps it to 00.
@@
for the original variable $b_{2}$ prior to the bias (5.10).
@@
6. Constructive training of shallow network
In this section, we follow the proof of Theorem 3.1 in order to determine a constructively trained shallow network, which does not invoke the gradient descent algorithm.
In our approach, all relevant operations are confined to a $Q\leq M$ dimensional linear subspace of ${\mathbb{R}}^{M}$. To this end, we calculate
$\overline{X_{0}^{red}}:=[\overline{x_{0,1}}\cdots\overline{x_{0,Q}}]\in{\mathbb{R}}^{M\times Q}\,.$ (6.1)
We determine the Penrose inverse
${\rm Pen}[\overline{X_{0}^{red}}]=\big{(}(\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\big{)}^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}$ (6.2)
and the orthoprojector
$P:=\overline{X_{0}^{red}}{\rm Pen}[\overline{X_{0}^{red}}]\in{\mathbb{R}}^{M\times M}\,.$ (6.3)
Subsequently, we determine $W_{1}=R\in O(M)$ which acts by rotating the ranges of $P$ and $P^{\perp}$ to be in alignment with the coordinate vectors $e_{j}:=(0,\dots 0,1,0,\dots,0)$. As $P$ is symmetric, this amounts to solving the diagonalization problem
$P=R^{T}P_{R}R\,.$ (6.4)
$P_{R}$ is diagonal, with $Q$ diagonal entries equaling $1$, and all other $M^{2}-Q$ matrix entries being 00. The orthogonal matrix $R$ diagonalizing $P$ also diagonalizes $P^{\perp}={\bf 1}-P$. We let
$W_{1}^{*}=R\in O(M)\,.$ (6.5)
Moreover, we set
$W_{2}^{*}=\widetilde{W}_{2}R^{T}$ (6.6)
where we calculate
$\widetilde{W}_{2}=Y((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}\,.$ (6.7)
To determine $B_{1}^{*}$, we split $b_{1}^{*}=P_{R}b_{1}^{*}+P^{\perp}_{R}b_{1}^{*}$, and set
$P_{R}b_{1}^{*}=\beta_{1}P_{R}u_{M}\,,$ (6.8)
recalling $u_{M}=(1,1,\dots,1)^{T}\in{\mathbb{R}}^{M}$ from (4.25). We choose $\beta_{1}\geq 0$ large enough to ensure that $RPX_{0}+P_{R}B_{1}^{*}$ is component-wise strictly positive. A sufficient condition is $\beta_{1}\geq 2\max_{j,i}|x_{0,j,i}|$. Moreover, we set
$P_{R}^{\perp}b_{1}^{*}=-\delta P^{\perp}_{R}u_{M}\,$ (6.9)
with $\delta$ as defined in (3.8). This defines $B_{1}^{*}=[b_{1}^{*}\cdots b_{1}^{*}]\in{\mathbb{R}}^{M\times N}$. Finally, we let
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\,.$ (6.10)
The set $W_{1}^{*},W_{2}^{*},b_{1}^{*},b_{2}^{*}$ defines the constructively trained shallow network. No use of the gradient descent algorithm is required.
@@ 6. Constructive training of shallow network
Here, we prove Theorem 3.3, and elucidate the geometric meaning of the constructively trained network obtained in Section 6. Given a test input $x\in{\mathbb{R}}^{M}$, the constructively trained network metrizes the linear subspace ${\rm range}(P)$ of the input space ${\mathbb{R}}^{M}$, and determines to which equivalence class of training inputs $Px$ is closest. To elucidate this interpretation, we observe that (6.7) implies that

### v3

@@ 2. Definition of the mathematical model
is invertible, and we can define the Penrose inverse [^1] Also called Moore–Penrose inverse or pseudoinverse and denoted $A^{+}$ instead of ${\rm Pen}[A]$. of $\overline{X_{0}^{red}}$,
${\rm Pen}[\overline{X_{0}^{red}}]:=\big{(}(\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}}\big{)}^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}.$ (2.24)
@@
[…]
@@
As the objective function, we consider the ${\mathcal{L}}^{2}$ Schatten class (or Hilbert-Schmidt) cost function
with
$\|A\|_{{\mathcal{L}}^{2}}\equiv\sqrt{{\rm Tr}(AA^{T})}\,,$ (2.28)
where $A^{T}$ is the transpose of the matrix $A$. Equivalently,
$\mathcal{C}[W_{j},b_{j}]$ $=\sqrt{\frac{1}{N}\sum_{j=1}^{Q}\sum_{i=1}^{N_{j}}|W_{2}\sigma(W_{1}x_{0,j,i}+b_{1})+b_{2}-y_{j}|^{2}_{{\mathbb{R}}^{Q}}}.$ (2.29)
The cost is a function of the weights and biases, while the training inputs in $X_{0}$ are considered to be a fixed set of parameters. Training this shallow network amounts to finding the optimizers $W_{j}^{**},b_{j}^{**}$ for the minimization problem
3. Statement of Main Results
All bounds derived here make use of the following quantity,
@@ 3. Statement of Main Results
which measures the relative size between $\overline{X_{0}^{red}}$ and $P\Delta X_{0}$, as $\left|{\rm Pen}[\overline{X_{0}^{red}}]P\Delta x_{0,j,i}\right|$ scales like the noise-to-signal ratio $\frac{|\Delta x|}{|x|}$ of training inputs. Note that it is scaling invariant under $X_{0}\rightarrow\lambda X_{0}$.
3.1. Upper bound on minimum of cost function for $M>Q$
Let $R\in O(M)$ denote an orthogonal matrix that diagonalizes $P$, defined in (2.25). That is, the orthoprojectors
$P_{R}:=RPR^{T}\;\;\;{\rm and}\;P^{\perp}_{R}:=RP^{\perp}R^{T}$ (3.2)
Letting
$u_{M}:=(1,1,\cdots,1)^{T}\in{\mathbb{R}}^{M}$ (3.3)
Then, we construct an upper bound on the cost function by use of the following weights and biases. We choose
$W_{1}^{*}=R$ (3.4)
and $b_{1}^{*}=P_{R}b_{1}^{*}+P_{R}^{\perp}b_{1}^{*}$ where
$\sigma(RPX_{0}+P_{R}B_{1}^{*})=RPX_{0}+P_{R}B_{1}^{*}\,.$ (3.6)
$\sigma(RP^{\perp}X_{0}+P_{R}^{\perp}B_{1}^{*})=0\,.$ (3.7)
Passing to the output layer, we require $W_{2}^{*}$ to solve
$W_{2}^{*}R\overline{X_{0}^{red}}=Y\,,$ (3.8)
which yields
$W_{2}^{*}=Y\;{\rm Pen}[\overline{X_{0}^{red}}]PR^{T}\,.$ (3.9)
Finally, we find that
$b_{2}^{*}=-W_{2}^{*}P_{R}b_{1}^{*}\,,$ (3.10)
that is, it reverts the translation by $P_{R}b_{1}^{*}$ in the previous layer.
@@ 3.1. Upper bound on minimum of cost function for $M>Q$
Therefore, the upper bound in (3.15) is scaling invariant, as it only depends on the signal-to-noise ratio of training input data, which is controlled by $\delta_{P}$.
@@ 3.2. Exact degenerate local minimum in the case $M=Q$
We explicitly determine an exact local degenerate minimum of the cost function in the case $Q=M$ where the input and output spaces have the same dimension; here, we show that the upper bound obtained above differs from the sharp value by a relative error of order $O(\delta_{P}^{2})$. We note that in this situation, the projector defined in (2.25) satisfies $P={\bf 1}_{Q\times Q}$ and $P^{\perp}=0$. Moreover, $\overline{X_{0}^{red}}$ is invertible, and therefore, ${\rm Pen}[\overline{X_{0}^{red}}]=(\overline{X_{0}^{red}})^{-1}$.
@@ 3.3. Geometric interpretation
The proof of Theorem 3.1 provides us with a particular set of weights $W_{1}^{*},W_{2}^{*}$ and biases $b_{1}^{*},b_{2}^{*}$ through an explicit construction, which yield the upper bound (3.15). We will refer to the shallow network trained with this choice of $W_{i}^{*},b_{i}^{*}$ as the constructively trained shallow network.
@@
and matching an input $x\in{\mathbb{R}}^{M}$ with a target output $y_{j^{*}}$ via the constructively trained shallow network is equivalent to the solution of the metric minimization problem
@@ 3.4. Dependence on truncation
We define the truncation map $\tau_{W_{1},b_{1}}$ as follows. It will play a major role in our explicit construction of local and global cost minimizers in deep neural networks in [CE23].
@@
Let $W_{1}\in GL(M)$, $b_{1}\in{\mathbb{R}}^{M}$. Then, the truncation map $\tau_{W_{1},b_{1}}:{\mathbb{R}}^{M\times N}\rightarrow{\mathbb{R}}^{M\times N}$ is defined by
@@
measures the signal-to-noise ratio of the truncated training input data.
@@
where $P_{R}$ has $Q$ diagonal entries of 111, and all other components are 0, while $P^{\perp}_{R}$ has $M-Q$ diagonal entries of 111, and all other components are 0. This does not characterize $R$ uniquely, as we may compose $R$ with an arbitrary $R^{\prime}\in O(M)$ that leaves the ranges of $P,P^{\perp}$ invariant. Diagonality of $P_{R},P_{R}^{\perp}$ implies that the activation function, which acts component-wise as the ramp function (2.5), satisfies
@@
is a matrix all of whose components are 0$\leq$ 0≤ 0. Therefore, $\sigma$ maps it to 0.
@@
for the original variable $b_{2}$ prior to the transformation (4.29).
@@
Here, we prove Theorem 3.3, and elucidate the geometric meaning of the constructively trained network obtained in Theorem 3.1. Given a test input $x\in{\mathbb{R}}^{M}$, the constructively trained network metrizes the linear subspace ${\rm range}(P)$ of the input space ${\mathbb{R}}^{M}$, and determines to which equivalence class of training inputs $Px$ is closest. To elucidate this interpretation, we observe that
$\widetilde{W}_{2}=Y((\overline{X_{0}^{red}})^{T}\overline{X_{0}^{red}})^{-1}(\overline{X_{0}^{red}})^{T}\in{\mathbb{R}}^{Q\times M}\,$ (6.1)
implies that